If the momentum of a body is increased by 100%the percentage increases in its kinetic energy is
1
Expert's answer
2018-10-22T11:29:09-0400

If the momentum of a body is increased by 100% the percentage, we get

p_2=2p_1

The kinetic energy

K=(mv^2)/2=p^2/2m

So

K_2=(p_2^2)/2m=(4p_1^2)/2m=4K_1

Thus, the kinetic energy of a body is increased by four times or by 300%.

Answer: 300%